OpenStudy (anonymous):

john and 2 friends are going out for pizza. They split one pizza and 3 drinks. The pizza costs 14. after using a 7 dollar coupon, they spend a total of 12.10 write an equation to model this situation find the cost for one drink

OpenStudy (anonymous):

14x+1.70y - c =Total Where x=the number of pizzas and y=the number of drinks and c is the coupon (assuming there is no tax).

zepdrix (zepdrix):

p=a pizza d=a drink p + 3d -7 = 12.10 <-- 1 pizza + 3 drinks - coupon = total bill 14 + 3d -7 = 12.10 3d + 7 = 12.10 3d = 5.10 d = 1.70 <-- cost for 1 drink.